### v1

@@ Acknowledgements.
Table 1 lists the properties of the spectroscopic sample. Lines used for atmospheric parameters, iron and titanium abundance estimation are listed in Table 2, while those used for chemical abundances are highlighted in Table 3. Table 4 shows the chemical abundances of the 3 RGB stars of M67 and the Sun. Finally, table 5 lists the values of for each spectral line computed as the mean abundance variation over 10 spectra, in response to changes in the atmospheric parameters.
@@
| HARPS1 | 231 | 29 | 115000 | [3790-6900] |
| SES2 | 400 | 64 | 55000 | [3860-8800] |
| FEROS3 | 339 | 161 | 48000 | [3500-9200] |
| UVES4 | 373 | 224 | 40000 | [3800-9400] |
| HARPS-N5 | 40 | 29 | 115000 | [3780-6910] |
| ESPaDOnS6 | 16 | 13 | 81000 | [3700-10500] |
1
High Accuracy Radial velocity Planet Searcher (HARPS, mayor03)
2
STELLA échelle Spectrograph (SES, strassmeier04; strassmeier10)
3
The Fiber-fed Extended Range Optical Spectrograph (FEROS, kaufer99)
4
Ultraviolet and Visual Echelle Spectrograph (UVES, dekker00)
5
High Accuracy Radial velocity Planet Searcher for the Northern hemisphere (HARPS-N, cosentino12)
6
Echelle SpectroPolarimetric Device for the Observation of Stars (ESPaDOnS, donati06)
@@
| X | -0.1 | -0.2 | +0.1 | +0.2 | -0.05 | -0.1 | +0.05 | +0.1 | -100 | -50 | +100 | +50 | -0.2 | -0.4 | +0.2 | +0.4 | |
| 4874.009 | Ti | 0.111 | 0.070 | 0.256 | 0.291 | 0.280 | 0.338 | 0.162 | 0.105 | 0.192 | 0.207 | 0.252 | 0.237 | 0.300 | 0.375 | 0.154 | 0.095 |
| 5105.530 | Cu | -0.002 | -0.007 | 0.009 | 0.017 | 0.048 | 0.092 | -0.034 | -0.074 | -0.074 | -0.037 | 0.095 | 0.046 | 0.033 | 0.063 | -0.017 | -0.041 |
| 5211.530 | Ti | 0.117 | 0.087 | 0.190 | 0.220 | 0.201 | 0.238 | 0.119 | 0.082 | 0.149 | 0.153 | 0.174 | 0.166 | 0.212 | 0.267 | 0.112 | 0.078 |
| 5218.195 | Cu | -0.012 | 0.001 | -0.006 | -0.010 | 0.055 | 0.125 | -0.072 | -0.126 | -0.088 | -0.043 | 0.062 | 0.027 | -0.004 | 0.034 | -0.018 | -0.014 |
| 5219.702 | Ti | 0.150 | 0.151 | 0.161 | 0.156 | 0.167 | 0.168 | 0.150 | 0.129 | 0.099 | 0.137 | 0.177 | 0.172 | 0.164 | 0.163 | 0.157 | 0.155 |
| 5220.065 | Cu | -0.057 | -0.162 | 0.173 | 0.249 | 0.227 | 0.321 | 0.009 | -0.165 | -0.009 | 0.036 | 0.316 | 0.183 | 0.121 | 0.124 | -0.013 | -0.075 |
| 5260.387 | Ca | -0.024 | -0.043 | 0.012 | 0.026 | 0.054 | 0.109 | -0.080 | -0.157 | -0.031 | -0.026 | 0.019 | 0.010 | -0.003 | -0.011 | -0.010 | -0.016 |
| 5381.022 | Ti | -0.223 | -0.239 | -0.138 | -0.129 | -0.135 | -0.125 | -0.178 | -0.200 | -0.165 | -0.159 | -0.143 | -0.148 | -0.129 | -0.113 | -0.201 | -0.239 |
| 5394.683 | Mn | -0.013 | -0.018 | 0.005 | 0.012 | 0.030 | 0.061 | -0.043 | -0.084 | -0.112 | -0.054 | 0.099 | 0.053 | -0.001 | -0.002 | 0.007 | 0.000 |
| 5418.768 | Ti | -0.039 | -0.072 | 0.043 | 0.078 | 0.052 | 0.093 | -0.033 | -0.073 | -0.016 | -0.005 | 0.030 | 0.020 | 0.088 | 0.137 | -0.061 | -0.119 |
| 5420.328 | Mn | -0.003 | -0.007 | 0.006 | 0.005 | 0.053 | 0.102 | -0.049 | -0.098 | -0.083 | -0.041 | 0.094 | 0.049 | 0.006 | 0.010 | -0.001 | -0.004 |
| 5557.063 | Al | -0.035 | -0.038 | -0.011 | -0.004 | 0.041 | 0.091 | -0.062 | -0.110 | -0.057 | -0.042 | 0.044 | 0.007 | -0.020 | -0.016 | -0.013 | -0.009 |
| 5588.749 | Ca | -0.005 | -0.006 | 0.005 | 0.014 | 0.060 | 0.120 | -0.066 | -0.127 | -0.066 | -0.036 | 0.067 | 0.032 | 0.142 | 0.304 | -0.130 | -0.227 |
| 5590.114 | Ca | 0.031 | 0.033 | 0.024 | 0.033 | 0.086 | 0.140 | -0.027 | -0.084 | -0.045 | -0.017 | 0.115 | 0.070 | 0.070 | 0.130 | -0.035 | -0.055 |
| 5645.613 | Si | 0.008 | 0.011 | 0.004 | 0.002 | 0.067 | 0.126 | -0.057 | -0.098 | -0.059 | -0.028 | 0.063 | 0.037 | 0.017 | 0.026 | -0.003 | -0.005 |
| 5665.555 | Si | 0.000 | 0.001 | -0.001 | 0.000 | 0.047 | 0.096 | -0.039 | -0.073 | -0.046 | -0.024 | 0.048 | 0.024 | 0.008 | 0.016 | -0.004 | -0.006 |
| 5684.484 | Si | -0.027 | 0.005 | 0.010 | 0.017 | 0.054 | 0.085 | -0.065 | -0.134 | -0.067 | -0.030 | 0.054 | 0.031 | 0.035 | 0.068 | -0.028 | -0.057 |
| 5688.205 | Na | -0.002 | 0.006 | -0.001 | 0.002 | 0.085 | 0.165 | -0.066 | -0.124 | -0.064 | -0.036 | 0.054 | 0.025 | 0.062 | 0.154 | -0.061 | -0.113 |
| 5700.170 | Cu | 0.005 | -0.031 | 0.008 | 0.028 | 0.127 | 0.250 | -0.153 | -0.306 | -0.103 | -0.065 | 0.123 | 0.059 | 0.019 | 0.023 | 0.013 | 0.005 |
| 5701.104 | Si | 0.007 | 0.006 | 0.002 | 0.001 | 0.051 | -0.012 | -0.032 | -0.072 | -0.034 | -0.014 | 0.046 | 0.025 | 0.011 | 0.019 | -0.002 | -0.006 |
| 5711.088 | Mg | 1.010 | 1.009 | 0.709 | 0.709 | 0.749 | 0.791 | 0.670 | 0.631 | 0.673 | 0.690 | 0.744 | 0.725 | 0.732 | 0.761 | 0.683 | 0.664 |
| 5867.562 | Ca | 0.018 | 0.021 | 0.007 | 0.003 | 0.057 | 0.093 | -0.018 | -0.042 | -0.020 | -0.008 | 0.061 | 0.035 | 0.010 | 0.013 | 0.016 | 0.024 |
| 6021.771 | Mn | 0.014 | 0.018 | 0.007 | 0.006 | 0.062 | 0.120 | -0.050 | -0.100 | -0.094 | -0.043 | 0.108 | 0.058 | 0.037 | 0.070 | -0.016 | -0.037 |
| 6154.226 | Na | 0.007 | 0.015 | -0.009 | -0.016 | 0.050 | 0.101 | -0.046 | -0.088 | -0.057 | -0.029 | 0.053 | 0.026 | 0.007 | 0.016 | -0.007 | -0.011 |
| 6160.747 | Na | 0.008 | 0.013 | -0.006 | -0.010 | 0.057 | 0.113 | -0.053 | -0.103 | -0.054 | -0.027 | 0.052 | 0.025 | 0.020 | 0.040 | -0.017 | -0.032 |
| 6166.439 | Ca | -0.002 | 0.031 | 0.008 | 0.016 | 0.048 | 0.107 | -0.033 | -0.094 | -0.035 | -0.031 | 0.067 | 0.033 | 0.031 | 0.064 | -0.026 | -0.022 |
| 6319.237 | Mg | 1.064 | 1.072 | 0.745 | 0.740 | 0.801 | 0.851 | 0.695 | 0.662 | 0.696 | 0.719 | 0.799 | 0.779 | 0.752 | 0.763 | 0.741 | 0.737 |
| 6371.371 | Si | -0.323 | -0.173 | -0.001 | 0.041 | -0.159 | 0.064 | -0.104 | -0.358 | 0.016 | -0.014 | -0.223 | -0.253 | -0.132 | 0.006 | -0.208 | -0.177 |
| 6455.598 | Ca | -0.006 | -0.013 | 0.016 | 0.027 | 0.073 | 0.137 | -0.054 | -0.133 | -0.054 | -0.024 | 0.072 | 0.037 | 0.016 | 0.029 | -0.013 | 0.003 |
| 6471.662 | Ca | 0.014 | 0.003 | 0.028 | 0.045 | 0.089 | 0.148 | -0.048 | -0.083 | -0.066 | -0.033 | 0.123 | 0.074 | 0.071 | 0.123 | -0.022 | -0.057 |
| 6499.650 | Ca | 0.032 | 0.046 | 0.017 | 0.022 | 0.071 | 0.124 | -0.031 | -0.081 | -0.037 | -0.016 | 0.089 | 0.059 | 0.046 | 0.084 | -0.000 | -0.018 |
| 6696.023 | Al | 0.001 | -0.001 | 0.000 | -0.005 | 0.073 | 0.138 | -0.066 | -0.126 | -0.076 | -0.039 | 0.082 | 0.041 | 0.009 | 0.017 | -0.005 | -0.011 |
| 6698.673 | Al | -0.003 | -0.004 | -0.001 | -0.007 | 0.054 | 0.114 | -0.063 | -0.110 | -0.073 | -0.037 | 0.066 | 0.029 | 0.004 | 0.008 | -0.009 | -0.013 |
| 6743.540 | S | -0.205 | -0.315 | 0.111 | 0.201 | -0.118 | 0.035 | 0.147 | -0.176 | -0.120 | 0.064 | -0.076 | 0.084 | 0.085 | -0.097 | -0.238 | -0.031 |
| 6757.150 | S | -0.200 | -0.173 | -0.101 | -0.081 | -0.094 | -0.049 | -0.063 | -0.300 | 0.016 | 0.008 | -0.166 | -0.133 | -0.083 | -0.014 | -0.086 | -0.094 |
@@ Appendix B New distance estimates for two Cepheids
The targets and have been estimated respectively at a distance of 5.64 kpc (andrievsky16) and of 10.04 kpc (trentin23).
andrievsky16 applied the method of the colour excess determination based on the use of the calibrating relation from friedman11. Here EW(DIB) is the equivalent width of the 6613 diffuse interstellar band (DIB), which is also seen in Cepheid spectra. Using this relation, they found that the value for is equal to 1.41. Combining the approximate relation , the ”absolute magnitude–pulsational period” relation of gieren98, and the mean V magnitude, they found a heliocentric distance of 5641 pc, and kpc. Given the , consistent with the rest of the sample, we used the Gaia parallax to obtain a heliocentric distance of kpc and kpc ( table, bailerjones21). Given this value so different from that obtained by the other authors, we used the W1-band PL relation, calibrated from Galactic CCs by wang18, obtaining a distance of kpc, consistent with the value that is obtained by the parallax method.
trentin23 found a heliocentric distance of kpc for using PW relations calibrated on Galactic CCs by ripepi20 and on LMC CCs by ripepi22. This measurement is in contrast with that of skowron19, who found 7.60.3 kpc using mid-IR PL relations. In a coherent manner as described in Sec. 2.1, we found a heliocentric distance of kpc using the Gaia DR3 parallax, which is consistent with that of skowron19 within 2.
@@ Appendix C Spectral synthesis details
Atmosphere model: we utilized MARCS12 (gustafsson2008grid).
@@
NLTE grids: we used amarsi2020galah grids.
@@
Initial chemical abundances: we used those from asplund09.
@@
PySME is designed to work in conjunction with the Vienna Atomic Line Database (VALD, ryabchikova15), which provides a comprehensive collection of atomic and molecular transition parameters of astronomical interest. In order to have good estimators we selected spectral lines with accurate atomic parameters. We updated the values for the lines used as estimators by referring to the Gaia-ESO database (heiter21). Gaia-ESO provides two quality parameters for each line: gfflag, which assesses the accuracy of , and synflag, which evaluates the blending resolution between closely spaced lines. These parameters are rated as ”y” (recommended use), ”n” (not recommended), or ”u” (undecided quality). For the line selection we need a relatively large number of estimators, possibly both neutral and ionized lines to be sensitive to the gravity variations. We selected Fe I, Fe II, Ti I, and Ti II lines with gfflag and synflag rated as ”y” for their high quality. This selection includes 21 Fe I lines, 2 Fe II lines, 6 Ti I lines, and 1 Ti II line. To expand the Fe II line list, we also included lines with gfflag rated as ”u”, adding 3 additional lines. Similarly, for the Ti II line list, we included lines with synflag rated as ”u”, adding 3 more lines. Fig. 21 displays an example of the fit of a Ti I line in a Cepheid spectrum. The complete line list is presented in Table 2.
We compared the atmospheric parameters derived from our method and the reference values for the Sun and 16 benchmark stars in Fig. 22.
We point out a limitation given by the atmosphere model used, since the atmospheric parameters are estimated at the edge of the grid for 22% of the spectra, as shown in Figure 23, where a ”staircase structure” artifact is in place. This implies that for the objects at the edge of the grid the surface gravity is in general overestimated, and/or the effective temperature is underestimated. As shown in Fig. 24, we compared our values of and with those of da2023oxygen to quantify this effect. We found an average difference of -0.20 dex in and 5 K in . While for the temperature the underestimation is completely negligible, for the surface gravity the average difference is comparable with the errors. For this reason, our final estimations of the surface gravity uncertainties take into account this effect.
@@
Furthermore, to provide a more quantitative analysis of the differences between the NLTE and LTE approaches, we made use of the large number of HR spectra collected by our group for 20 calibrating Cepheids (da2022new). For these objects, the HR spectra cover the entire pulsation cycle, making them an excellent benchmark for validating spectroscopic approaches. The reason is twofold. a) Classical Cepheids with moderate to large amplitudes show variations along the pulsation cycle of approximately 1000 K in , about half a dex in , and a factor of two in (from 2 to 4 ). However, the chemical abundances for each object are the same. This makes them a robust laboratory for constraining the physical assumptions involved when transitioning from LTE to NLTE analyses. b) Classical Cepheids trace a loop in both the color-magnitude diagram and the spectroscopic colour-magnitude diagram (or Kiel diagram; vs. ; see kudritzki20, kudritzki24). As a consequence, during the pulsation cycle a Cepheid can reach, at a fixed surface gravity, two different effective temperatures: one on the rising branch and one on the declining branch. This provides a unique opportunity to test the impact of on NLTE analysis while keeping both surface gravity and chemical composition fixed.
We focused our analysis on three Cepheids for which we have very detailed phase coverage: , (fundamental mode; Fig. 26), (fundamental mode; Fig. 27), and (first overtone; Fig. 28). The bottom four panels of these figures show the differences in atmospheric parameters and iron abundances between the NLTE and LTE analyses. These panels reveal overall consistency between the two approaches, as no significant trends or systematic offsets are observed. The same conclusion holds when comparing the LTE results from da2022new with our current NLTE analysis (top four panels of the same figures).

### v2

@@ Appendix A Tables
Table 1 lists the properties of the spectroscopic sample. Lines used for atmospheric parameters, iron and titanium abundance estimation are listed in Table 11, while those used for chemical abundances are highlighted in Table 12. Table 4 shows the chemical abundances of the 3 giant stars of M67 and the Sun. Furthermore, Table 13 lists the values of for each spectral line computed as the mean abundance variation over 10 spectra, in response to changes in the atmospheric parameters. Finally, Table 14 lists all the kinematic and chemical information for each Cepheid, with the flag ”” which determines whether the Cepheid has been added to the original dS23 sample, while the information and the results for each spectrum are highlighted in Table 15.
@@
| HARPSaaaaHigh Accuracy Radial velocity Planet Searcher (HARPS, Mayor et al. 2003). | 231 | 29 | 115000 | [3790-6900] |
| SESbbbbSTELLA échelle Spectrograph (SES, Strassmeier et al. 2004, 2010). | 400 | 64 | 55000 | [3860-8800] |
| FEROSccccThe Fiber-fed Extended Range Optical Spectrograph (FEROS, Kaufer et al. 1999). | 339 | 161 | 48000 | [3500-9200] |
| UVESddddUltraviolet and Visual Echelle Spectrograph (UVES, Dekker et al. 2000). | 373 | 224 | 40000 | [3800-9400] |
| HARPS-NeeeeHigh Accuracy Radial velocity Planet Searcher for the Northern hemisphere (HARPS-N, Cosentino et al. 2012). | 40 | 29 | 115000 | [3780-6910] |
| ESPaDOnSffffEchelle SpectroPolarimetric Device for the Observation of Stars (ESPaDOnS, Donati et al. 2006). | 16 | 13 | 81000 | [3700-10500] |
@@
| 0.1 | 0.2 | +0.1 | +0.2 | 0.05 | 0.1 | +0.05 | +0.1 | 100 | 50 | +100 | +50 | 0.2 | 0.4 | +0.2 | +0.4 | ||
| 4874.01 | Ti | 0.111 | 0.070 | 0.256 | 0.291 | 0.280 | 0.338 | 0.162 | 0.105 | 0.192 | 0.207 | 0.252 | 0.237 | 0.300 | 0.375 | 0.154 | 0.095 |
| 5105.53 | Cu | 0.002 | 0.007 | 0.009 | 0.017 | 0.048 | 0.092 | 0.034 | 0.074 | 0.074 | 0.037 | 0.095 | 0.046 | 0.033 | 0.063 | 0.017 | 0.041 |
| 5211.53 | Ti | 0.117 | 0.087 | 0.190 | 0.220 | 0.201 | 0.238 | 0.119 | 0.082 | 0.149 | 0.153 | 0.174 | 0.166 | 0.212 | 0.267 | 0.112 | 0.078 |
| 5218.20 | Cu | 0.012 | 0.001 | 0.006 | 0.010 | 0.055 | 0.125 | 0.072 | 0.126 | 0.088 | 0.043 | 0.062 | 0.027 | 0.004 | 0.034 | 0.018 | 0.014 |
| 5219.70 | Ti | 0.150 | 0.151 | 0.161 | 0.156 | 0.167 | 0.168 | 0.150 | 0.129 | 0.099 | 0.137 | 0.177 | 0.172 | 0.164 | 0.163 | 0.157 | 0.155 |
| 5220.07 | Cu | 0.057 | 0.162 | 0.173 | 0.249 | 0.227 | 0.321 | 0.009 | 0.165 | 0.009 | 0.036 | 0.316 | 0.183 | 0.121 | 0.124 | 0.013 | 0.075 |
| 5260.39 | Ca | 0.024 | 0.043 | 0.012 | 0.026 | 0.054 | 0.109 | 0.080 | 0.157 | 0.031 | 0.026 | 0.019 | 0.010 | 0.003 | 0.011 | 0.010 | 0.016 |
| 5381.02 | Ti | 0.223 | 0.239 | 0.138 | 0.129 | 0.135 | 0.125 | 0.178 | 0.200 | 0.165 | 0.159 | 0.143 | 0.148 | 0.129 | 0.113 | 0.201 | 0.239 |
| parallax over error | pmra | pmdec | RUWE | N25 | dS23 | d | … | |||||||
| [deg] | [deg] | [mas] | [mas] | [] | [] | [pc] | [pc] | |||||||
| AA Gem | 3430067092837622272 | 91.646 | 26.329 | 0.275 | 0.018 | 15.548 | 0.217 | 0.670 | 1.249 | n | y | 3245 | 184 | … |
| AB Vel | 5354898841641168768 | 154.967 | 56.310 | 0.229 | 0.012 | 18.876 | 5.640 | 2.616 | 0.907 | y | y | 4430 | 235 | … |
| AC Mon | 3050050207554658048 | 105.249 | 8.709 | 0.355 | 0.019 | 19.010 | 2.237 | 1.964 | 1.379 | y | y | 2631 | 126 | … |
| AD Cru | 6057514092119497472 | 183.249 | 62.097 | 0.293 | 0.013 | 22.121 | 5.338 | 0.473 | 1.017 | y | y | 3220 | 149 | … |
| AD Pup | 5614312705966204288 | 117.016 | 25.578 | 0.233 | 0.017 | 14.088 | 2.358 | 3.054 | 1.362 | y | y | 3995 | 276 | … |
| AE Tau | 3441145084801214464 | 83.653 | 26.202 | 0.274 | 0.020 | 14.044 | 0.071 | 1.252 | 1.050 | y | y | 3347 | 231 | … |
| AE Vel | 5309174967720762496 | 144.214 | 53.033 | 0.353 | 0.012 | 29.138 | 5.137 | 4.069 | 0.969 | y | y | 2726 | 99 | … |
| AG Cru | 6059635702888301952 | 190.358 | 59.794 | 0.742 | 0.020 | 37.358 | 4.422 | 1.847 | 1.016 | y | y | 1323 | 32 | … |
| spectrum ID | MJD | instrument | Star | Gaia ID | … | |||||||
| [K] | [K] | [dex] | [dex] | [dex] | [dex] | |||||||
| ADP.2014-09-26T16_50_57.353 | 56213.9879650 | HARPS | V496 Aql | 4204653587029046400 | 5664 | 140 | 1.02 | 0.15 | 0.12 | 0.08 | 3.55 | … |
| ADP.2014-09-26T16_54_13.337 | 56213.9822101 | HARPS | Y Oph | 4175017625462647168 | 5662 | 150 | 1.01 | 0.15 | 0.08 | 0.05 | 4.06 | … |
| ADP.2014-09-26T16_54_45.020 | 56239.9935618 | HARPS | V496 Aql | 4204653587029046400 | 5765 | 150 | 1.01 | 0.15 | 0.13 | 0.05 | 3.90 | … |
| ADP.2014-09-26T16_54_49.173 | 56212.9902382 | HARPS | Y Sgr | 4096107909387492992 | 5755 | 100 | 1.03 | 0.10 | 0.12 | 0.08 | 3.720 | … |
| ADP.2014-10-01T10_19_05.720 | 53034.1506477 | HARPS | Gem | 3366754155291545344 | 5528 | 90 | 1.25 | 0.10 | 0.05 | 0.08 | 3.08 | … |
| ADP.2014-10-01T10_19_06.007 | 53149.2916913 | HARPS | Y Sgr | 4096107909387492992 | 5564 | 130 | 1.43 | 0.15 | 0.00 | 0.05 | 3.96 | … |
| ADP.2014-10-01T10_19_06.560 | 53021.1981818 | HARPS | Gem | 3366754155291545344 | 5842 | 130 | 1.01 | 0.15 | 0.04 | 0.08 | 3.58 | … |
| ADP.2014-10-01T10_19_12.130 | 53025.1780493 | HARPS | Dor | 4757601523650165120 | 5270 | 100 | 0.99 | 0.10 | 0.11 | 0.08 | 3.21 | … |
@@ Appendix B New distance estimates for two Cepheids
The targets - and - have been estimated respectively at a distance of 5.64 kpc (Andrievsky et al. 2016) and of 10.04 kpc (Trentin et al. 2023).
Andrievsky et al. (2016) applied the method of the colour excess determination based on the use of the calibrating relation from Friedman et al. (2011). Here EW(DIB) is the equivalent width of the 6613 diffuse interstellar band (DIB), which is also seen in Cepheid spectra. Using this relation, they found that the value for - is equal to 1.41. Combining the approximate relation , the ”absolute magnitude–pulsational period” relation of Gieren et al. (1998), and the mean V magnitude, they found a heliocentric distance of 5641 pc, and kpc. Given the , we used the Gaia parallax to obtain a heliocentric distance of kpc and kpc ( table, Bailer-Jones et al. 2021). Given this value so different from that obtained by the other authors, we used the W1-band PL relation, calibrated from Galactic CCs by Wang et al. (2018), obtaining a distance of kpc, consistent with the value that is obtained by the parallax method.
Trentin et al. (2023) found a heliocentric distance of kpc for - using PW relations calibrated on Galactic CCs by Ripepi et al. (2020) and on LMC CCs by Ripepi et al. (2022). This measurement is in contrast with that of Skowron et al. (2019), who found kpc using mid-IR PL relations. In a coherent manner as described in Sec. 2.1, we found a heliocentric distance of kpc using the Gaia DR3 parallax, which is consistent with that of Skowron et al. (2019) within 2.
@@ Appendix C Spectral synthesis details
Atmosphere model: we utilized MARCS12 (Gustafsson et al. 2008).
@@
NLTE grids: we used Amarsi et al. (2020) grids.
@@
Initial chemical abundances: we used those from Asplund et al. (2009).
@@
PySME is designed to work in conjunction with the Vienna Atomic Line Database (VALD, Ryabchikova et al. 2015), which provides a comprehensive collection of atomic and molecular transition parameters of astronomical interest. In order to have good estimators we selected spectral lines with accurate atomic parameters. We updated the values for the lines used as estimators by referring to the Gaia-ESO database (Heiter et al. 2021). Gaia-ESO provides two quality parameters for each line: gfflag, which assesses the accuracy of , and synflag, which evaluates the blending resolution between closely spaced lines. These parameters are rated as ”y” (recommended use), ”n” (not recommended), or ”u” (undecided quality). For the line selection we need a relatively large number of estimators, possibly both neutral and ionized lines to be sensitive to the gravity variations. We selected Fe I, Fe II, Ti I, and Ti II lines with gfflag and synflag rated as ”y” for their high quality. This selection includes 21 Fe I lines, 2 Fe II lines, 6 Ti I lines, and 1 Ti II line. To expand the Fe II line list, we also included lines with gfflag rated as ”u”, adding 3 additional lines. Similarly, for the Ti II line list, we included lines with synflag rated as ”u”, adding 3 more lines. Fig. 21 displays an example of the fit of a Ti I line in a Cepheid spectrum. The complete line list is presented in Table 11. We compared the atmospheric parameters derived from our method and the reference values for the Sun and 16 benchmark stars in Fig. 22.
We point out a limitation given by the atmosphere model used, since the atmospheric parameters are estimated at the edge of the grid for 22% of the spectra, as shown in Figure 23, where a ”staircase structure” artifact is in place. This implies that for the objects at the edge of the grid the surface gravity is in general overestimated, and/or the effective temperature is underestimated. As shown in Fig. 24, we compared our values of and with those of dS23 to quantify this effect. We found an average difference of dex in and 5 K in . While for the temperature the underestimation is completely negligible, for the surface gravity the average difference is comparable with the errors. For this reason, our final estimations of the surface gravity uncertainties take into account this effect.
@@
Furthermore, to provide a more quantitative analysis of the differences between the NLTE and LTE approaches, we made use of the large number of HR spectra collected by our group for 20 calibrating Cepheids (da Silva et al. 2022). For these objects, the HR spectra cover the entire pulsation cycle, making them an excellent benchmark for validating spectroscopic approaches. The reason is twofold. a) Classical Cepheids with moderate to large amplitudes show variations along the pulsation cycle of approximately 1000 K in , about half a dex in , and a factor of two in (from 2 to 4 ). However, the chemical abundances for each object are the same. This makes them a robust laboratory for constraining the physical assumptions involved when transitioning from LTE to NLTE analyses. b) Classical Cepheids trace a loop in both the colour-magnitude diagram and the spectroscopic colour-magnitude diagram (or Kiel diagram; vs. ; see Kudritzki et al. 2020, 2024). As a consequence, during the pulsation cycle a Cepheid can reach, at a fixed surface gravity, two different effective temperatures: one on the rising branch and one on the declining branch. This provides a unique opportunity to test the impact of on NLTE analysis while keeping both surface gravity and chemical composition fixed.
We focused our analysis on three Cepheids for which we have very detailed phase coverage: , (fundamental mode; Fig. 26), (fundamental mode; Fig. 27), and (first overtone; Fig. 28). The bottom four panels of these figures show the differences in atmospheric parameters and iron abundances between the NLTE and LTE analyses. These panels reveal overall consistency between the two approaches, as no significant trends or systematic offsets are observed. The same conclusion holds when comparing the LTE results from da Silva et al. (2022) with our current NLTE analysis (top four panels of the same figures).